QUESTION 13 Amount ($b) 4.5 5.5 190 60 150 410 Risk Weight 0% 0% 5% 100% 100% Asset Risk Weighted Assets ES funds T-notes and CG bonds Home loans (LVR < 80%) Home loans (LVR > 85%) Business loans Total Calculate the risk weighted value of these assets to one decimal. Write $1 billion as 1QUESTION 14 Risk Weight 0% 0% 75% 100% 100% Risk Weighted Assets Amount ($b) 4.5 5.5 190 60 150 410 Asset ES funds T-notes and CG bonds Home loans (LVR < 80%) Home loans (LVR > 85%) Business loans Total The bank has $20b equity. Calculate the equity to risk weighted asset ratio (in 96, 196-1) You can use (your hopefully correct!) answer from before for the RWA (risk weighted assets)
Answer #1

Answer 13

ES FUND 4.5/410*0=0
T-NOTE AND CG BOND 5.5/410*0=0
HOME LOAN 190/410*.75=.3
HOME LOAN 60/410*.100=.1
BUSINESS LOAN 150/410*.100=.3
TOTAL =.7*410= 287

14.

risk weighted assets = $287

equity =$20

calculating equity to risk weighted asset ratio.

= equity/ risk weighted assets

20/287=$.06968